# Bayesian Speech synthesizers Can Learn from Multiple Teachers

### v1

@@ 5.2 Evaluation Settings
For subjective evaluation, MOS and SMOS scores are obtained via a crowd-sourcing platform. MOS for evaluating overall speech quality, and SMOS for measuring speaker similarity between the generated audio and the prompt. MOS and SMOS is evaluated following the detailed procedure described in Appendix  F.
@@
All evaluation details could be found in Appendix C.5.
@@ 6.1 Main Results
Table 1: Comparison of MOS and SMOS for different systems. BELLE-stream is fixed to a single speaker timbre and does not use an audio prompt, hence SMOS is not reported.
System Ground Truth MaskGCT F5-TTS MELLE BELLE BELLE-stream
MOS 4.20 4.12 4.25 4.02 4.21 4.06
SMOS 3.81 3.89 3.95 3.80 4.13
Table 2: Comparison of WER (%) and speaker similarity metrics for BELLE and baselines. Entries marked with † are trained using the dataset described in Sec. 5.1, marked with * correspond to the streaming TTS setting. BELLE-stream only takes text as input and does not use audio prompt, different from Cross-Sentence setting, so SIM metrics are omitted.
@@
MELLE † 2.04 2.59 0.526 0.488 3.30 3.83 0.652 0.606
SMLLE(Sun et al., 2025) * - - - - 5.14 6.37 0.516 0.489
IST-LM(Yang et al., 2024) * - 3.60 - - - 4.53 - 0.653
BELLE † 1.63 2.13 0.549 0.519 2.45 2.99 0.679 0.641
BELLE-stream* - - - - 2.70 3.97 - -
From Table 1, BELLE, F5-TTS, and Ground Truth achieve MOS scores of around 4.2, indicating that the audio produced by both F5-TTS and BELLE is close to that of natural human speech. F5-TTS attains a slightly higher MOS, which can be attributed to its lower background noise — human raters tend to prefer cleaner audio. In contrast, the real speech in Ground Truth inherently contains a certain level of noise, and BELLE replicates this natural background characteristic, leading to MOS scores that are very close to the Ground Truth. Regarding SMOS, it is worth noting that in the LibriSpeech dataset, different utterances from the same speaker sometimes exhibit perceptible timbre variation, which results in a relatively lower SMOS score for Ground Truth. BELLE achieves the highest speaker similarity among all tested systems. Notably, while both F5‑TTS and MaskGCT were trained on datasets comprising 50,000 hours of English speech, BELLE was trained on less than 5,000 hours of data, yet it attained competitive, and in some instances superior performance. These results underscore the efficacy of the proposed Bayesian sampling strategy and the multi‑teacher learning framework.
For the objective metrics in Table 2, BELLE attains the best WER and SIM scores under the Continuation setting. Under the Cross-sentence setting, MaskGCT achieves slightly higher scores; however, in our reproduction, MaskGCT’s WER-C is only 3.86% under our evaluation protocol. This suggests that BELLE remains highly competitive in the Cross-sentence scenario. Additionally, BELLE outperforms MELLE across both subjective and objective metrics, including MOS, SMOS, WER, and SIM. These results demonstrate that Bayesian sampling delivers superior performance compared to Gaussian sampling.
Since BELLE-stream is finetuned to a fixed speaker timbre, it does not accept an audio prompt; therefore, it is excluded from SMOS scoring in the subjective evaluation and SIM computation in the objective evaluation, and its results are reported only under the Cross-sentence setting. BELLE-stream only takes the target text to be synthesized as input. From Table 1, streaming generation does not cause significant degradation in speech naturalness, achieving higher MOS scores than non‑streaming MELLE. From Table 2, BELLE-stream achieves the lowest WER among streaming TTS systems, with no appreciable increase in WER compared to BELLE. In our tests, BELLE-stream achieved a real-time factor (RTF)—defined as the ratio of audio generation time to the length of the generated audio—of 0.55 when producing a 10-second utterance. Given that BELLE-stream operates with 0.8-second chunks, this implies a first-packet latency (FPL) of only 440ms, demonstrating its effective balance between performance and latency.
@@ 6.2 Diversity Analysis
The diversity evaluation is conducted using the evaluation method showed in Sec. 5.2. In the NIG prior, the parameter $\beta_{t}$ denotes the scale of the inverse-gamma distribution over the variance $\sigma_{t}^{2}$. Increasing $\beta_{t}$ raises the expected variance, thereby promoting greater diversity in the sampled outcomes (See Eqn. equation 7). From Table 3, it’s observed that BELLE ($\beta*2$) exhibits the largest distances across all three metrics, confirming that increasing the sampling parameter $\beta$ effectively enhances generation diversity. Default BELLE achieves moderate diversity, slightly higher than MELLE, suggesting that Bayesian sampling yields superior diversity compared to Gaussian sampling. In contrast, F5-TTS produces consistently small distances, indicating that its outputs are more deterministic and less diverse across repeated sampling. These findings highlight that, beyond intelligibility and speaker similarity, BELLE offers a controllable trade-off between stability and diversity at inference time, with the $\beta$ parameter serving as a practical knob to adjust sample variability, analogous to the role of the temperature parameter in the sampling process of token-based language models.
6.3 The effect of multi teachers
Due to limited computational resources, the analysis experiments are conducted on a smaller-scale dataset. The 1-teacher setting trains solely on the original Librispeech data without synthesized augmentation. The 3-teacher setting incorporates additional synthesized speech from two TTS models, while the 7-teacher setting augments the training data with speech from all six TTS models described earlier. The data-aug configuration serves as a baseline using conventional data augmentation, randomly sampling from all available data sources during training. Details about the small dataset and model training configuration could be found in Appendix C.4.
As shown in Table 4, the performance of BELLE consistently improves as the number of teachers increases. Furthermore, the possibility that the performance gain is solely due to the increased amount of training data is excluded. In the data-aug setting, the total training data is larger than that of the 3-teacher configuration; however, its performance is inferior to 3-teacher on multiple metrics. This degradation may be attributed to the conventional data augmentation strategy, in which the same text can correspond to acoustically different speeches, potentially causing confusion for the model. In contrast, under our proposed multi-teacher training scheme, increasing the number of teachers leads to consistent performance gains, thereby demonstrating the effectiveness of the multi teacher training strategy.
Table 4: Comparison of WER (%) and speaker similarity metrics for BELLE under different conditions trained on a smaller-scale dataset. 1-teacher indicates training with only the original Librispeech dataset; 3-teacher additionally includes XTTS-v2 and MaskGCT synthesized data; 7-teacher further includes synthesized data from six TTS models. data-aug randomly samples from combined all data sources of 7 teachers.
@@ 6.3 The effect of multi teachers
BELLE 1-teacher 3.10 3.92 0.344 0.303 6.11 7.34 0.374 0.330
BELLE 3-teacher 2.04 2.66 0.398 0.362 4.12 4.84 0.433 0.395
BELLE 7-teacher 1.96 2.57 0.444 0.409 4.17 4.73 0.485 0.449
@@
Table 5: Ablation study examining the effects of the sampling module and flux loss on BELLE 3-teacher. ✓ indicates enabled, ✗ indicates disabled.
Sampling Flux Loss Continuation Cross-Sentence
WER-C WER-H SIM-r SIM-o WER-C WER-H SIM-r SIM-o
4.30 5.00 0.382 0.344 12.04 13.07 0.392 0.351
2.50 3.25 0.388 0.347 5.61 6.37 0.417 0.372
2.04 2.66 0.398 0.362 4.12 4.84 0.433 0.395
6.4 The effects of the sampling module and flux loss
An ablation study is conducted to investigate how the sampling module and flux loss affect the performance of the BELLE model. Due to limited computational resources, the analysis experiments are also conducted on the smaller-scale dataset. Our experiments are based on the BELLE 3-teacher setting. As demonstrated in Table 5, the sampling module plays a more crucial role than flux loss. Removing the flux loss results in only a slight performance degradation, whereas removing the sampling module leads to a severe performance drop. This indicates that Bayesian sampling plays a crucial role in the effectiveness of BELLE.
@@ 7 Conclusion
In this paper, we propose BELLE, a novel continuous-valued AR TTS model integrating Bayesian EDL for sampling mel-spectrogram frames. Our approach addresses the underexplored area of effective sampling methods for continuous-valued AR TTS. Additionally, a multi-teacher knowledge distillation framework is introduced, substantially improving TTS synthesis quality using synthesized data from publicly available models. Experimental results highlight the efficacy of Bayesian evidential sampling and multi-teacher distillation in achieving competitive speech naturalness, speaker similarity, and diversity, rivaling models trained on much larger real-data corpora. Remarkably, our approach attains these results using only one-tenth of the data, much of it synthetic, while enabling high-quality, low-latency TTS suitable for both offline and streaming applications.
Reproducibility Statement
To ensure reproducibility, we provide detailed model configurations, training parameters, and training data specifications in Sec. 5 and Appendix C. Upon acceptance, we will release all code, checkpoints, and synthesized speech data, enabling the community to reliably reproduce our work and further explore its potential.
Ethics statement
@@ Appendix A Limitations
Although BELLE demonstrates strong performance, there are two main limitations. First, the current study focuses solely on training and evaluation with English speech data, without validating multilingual generalization; extending BELLE to diverse languages remains an important direction for future work. Second, as with most AR TTS systems, BELLE’s RTF is still higher than that of recent flow-matching-based models. Further research on predicting multiple mel-spectrogram frames per step could substantially reduce RTF and enable lower-latency streaming TTS.
Appendix B Use of Large Language Models
In the preparation of this work, large language models (LLMs) were employed to refine the writing for clarity and conciseness. During the experimental phase, LLMs also provided code modification suggestions that assisted in implementing and debugging parts of the system. All conceptual designs, experimental settings, and core contributions remain the authors’ original work.
Appendix C Detailed Experimental Setup
C.1 Training Data

### v2

@@ 5.2 Evaluation Settings
For subjective evaluation, MOS and SMOS scores are obtained via a crowd-sourcing platform. MOS for evaluating overall speech quality, and SMOS for measuring speaker similarity between the generated audio and the prompt. MOS and SMOS is evaluated following the detailed procedure described in Appendix  H.
@@
All evaluation details could be found in Appendix D.5.
@@ 6.1 Main Results
Table 1: Comparison of MOS and SMOS for different systems. Results are reported as mean $\pm$ standard deviation calculated across all ratings for each system.
System Ground Truth MaskGCT F5-TTS MELLE BELLE
MOS $4.20\pm 0.89$ $4.12\pm 0.85$ $\mathbf{4.25}\pm 0.81$ $4.02\pm 0.96$ $4.21\pm 0.75$
SMOS $3.81\pm 0.86$ $3.89\pm 0.90$ $3.95\pm 0.91$ $3.80\pm 0.90$ $\mathbf{4.13}\pm 0.78$
Table 2: Comparison of WER (%) and speaker similarity metrics. GT indicates training with only the original LibriSpeech dataset to validate BELLE’s effectiveness without synthetic data; GT + 6 TTS further includes synthesized data from all six TTS models.
@@
VALL-E (Wang et al., 2023) - 3.8 0.508 - - 5.9 0.580 -
MELLE GT 2.15 2.78 0.491 0.452 7.78 8.76 0.610 0.561
BELLE GT 1.79 2.39 0.492 0.457 3.81 5.01 0.620 0.576
MELLE GT + 6 TTS 2.04 2.59 0.526 0.488 3.30 3.83 0.652 0.606
BELLE GT + 6 TTS 1.63 2.13 0.549 0.519 2.45 2.99 0.679 0.641
Subjective Evaluation (MOS & SMOS).
From Table 1, BELLE, F5-TTS, and Ground Truth all achieve MOS scores around 4.2, suggesting that the generated audio is comparable to natural human speech. Notably, BELLE achieves this competitive performance using less than 5,000 hours of data, compared to $\sim$50,000 hours required by baselines like F5-TTS. While F5-TTS attains a marginally higher mean score (4.25)—likely due to its tendency to produce cleaner audio by removing background noise—BELLE demonstrates superior stability, achieving the lowest standard deviation ($\pm$ 0.75) among all systems, including Ground Truth ($\pm$ 0.89). This indicates that BELLE faithfully replicates the acoustic characteristics of the reference audio without the quality fluctuations often observed in other models. Regarding SMOS, despite the relatively lower Ground Truth score caused by inherent timbre variations in the LibriSpeech dataset, BELLE achieves the highest mean speaker similarity (4.13) with the smallest standard deviation ($\pm$ 0.78), significantly outperforming baselines like F5-TTS ($\pm$ 0.91). This confirms that BELLE captures target timbres more reliably and robustly.
Objective Evaluation.
In terms of objective metrics (Table 2), BELLE achieves state-of-the-art performance, surpassing all baseline systems in both Continuation and Cross-Sentence settings. Specifically, BELLE attains the lowest word error rates (WER-C 1.63% and 2.45%) and the highest speaker similarity scores, significantly outperforming strong baselines such as RALL-E and CLAM-TTS. Crucially, under identical training conditions, BELLE consistently outperforms MELLE across all metrics. For instance, in the Cross-Sentence setting, BELLE reduces WER-H by nearly 22% (from 3.83% to 2.99%) and improves SIM-r from 0.652 to 0.679. These results provide compelling empirical evidence that the proposed Bayesian evidential sampling framework delivers superior robustness and fidelity compared to standard Gaussian sampling.
The Role of Synthetic Data
To understand the source of BELLE’s efficiency and robustness, we analyze the distinct contributions of the Bayesian framework and the synthetic data strategy. First, this effectiveness is intrinsic to the framework, not reliant on synthetic data. As shown in Table 2 (GT setting), BELLE trained solely on human speech drastically outperforms MELLE (e.g., 3.81% vs. 7.78% WER-C). This confirms that the EDL formulation effectively captures uncertainty even from standard corpora by exploiting phonemic recurrence. Second, the synthetic data serves a statistical purpose, not an imitative one. Since standard datasets lack the “one-to-many” variations required for variance estimation, we use synthetic samples to characterize the distributional envelope. By aggregating diverse outputs from multiple teachers, our framework learns robust variance properties without overfitting to the artifacts or errors of any single teacher, keeping the core generation quality anchored in the original human recordings.
@@ 6.2 Diversity Analysis
The diversity evaluation is conducted using the evaluation method showed in Sec. 5.2. In the NIG prior, the parameter $\beta_{t}$ denotes the scale of the inverse-gamma distribution over the variance $\sigma_{t}^{2}$. Increasing $\beta_{t}$ raises the expected variance, thereby promoting greater diversity in the sampled outcomes (See Eqn. (10)). From Table 3, it’s observed that BELLE ($\beta*2$) exhibits the largest distances across all three metrics, confirming that increasing the sampling parameter $\beta$ effectively enhances generation diversity. Default BELLE achieves moderate diversity, slightly higher than MELLE, suggesting that Bayesian sampling yields superior diversity compared to Gaussian sampling. In contrast, F5-TTS produces consistently small distances, indicating that its outputs are more deterministic and less diverse across repeated sampling. These findings highlight that, beyond intelligibility and speaker similarity, BELLE offers a controllable trade-off between stability and diversity at inference time, with the $\beta$ parameter serving as a practical knob to adjust sample variability, analogous to the role of the temperature parameter in the sampling process of token-based language models.
6.3 The Effect of Teacher Counts
Table 4: Comparison of WER (%) and speaker similarity metrics for BELLE under different training configurations on a smaller-scale dataset. GT indicates training with only the original LibriSpeech dataset; GT + 2 TTS additionally includes synthesized data from XTTS-v2 and MaskGCT; GT + 6 TTS further includes synthesized data from all six TTS models. data-aug randomly samples from the combined pool of all data sources (GT and 6 TTS) without the evidential “one-to-many” alignment.
@@ 6.3 The Effect of Teacher Counts
BELLE GT 3.10 3.92 0.344 0.303 6.11 7.34 0.374 0.330
BELLE GT + 2 TTS 2.04 2.66 0.398 0.362 4.12 4.84 0.433 0.395
BELLE GT + 6 TTS 1.96 2.57 0.444 0.409 4.17 4.73 0.485 0.449
@@
Due to limited computational resources, the analysis experiments are conducted on a smaller-scale dataset. The GT setting trains solely on the original LibriSpeech data (Ground Truth) without synthesized augmentation. The GT + 2 TTS setting incorporates additional synthesized speech from two TTS models, while the GT + 6 TTS setting augments the training data with speech from all six TTS models described earlier. The data-aug configuration serves as a baseline using conventional data augmentation, randomly sampling from all available data sources during training. Details about the small dataset and model training configuration can be found in Appendix D.4.
As shown in Table 4, the performance of BELLE consistently improves as the diversity of data sources increases. Furthermore, we exclude the possibility that the performance gain is solely due to the increased volume of training data. In the data-aug setting, the total training data volume is strictly larger than that of the GT + 2 TTS configuration; however, its performance is inferior to GT + 2 TTS on multiple metrics. This degradation may be attributed to the conventional data augmentation strategy, in which the same text can correspond to acoustically different speeches in a random manner, potentially causing confusion for the model. In contrast, under our proposed “one-to-many” training scheme, expanding from GT to GT + 6 TTS leads to consistent performance gains, thereby demonstrating the effectiveness of the evidential training strategy.
Notably, our synthetic multi-sample speech is introduced to provide multiple realizations of the same text—required by the EDL-based Bayesian formulation for uncertainty estimation—rather than to replace human recordings or to imitate any specific teacher model. The core speech generation capability is still grounded in the real human speech (GT), while the synthetic samples act only as statistical observations to characterize conditional variance.
6.4 Results of BELLE-stream
Table 5: Comparison of WER (%) and speaker similarity metrics in the streaming setting.
System Cross-Sentence
WER-C WER-H SIM-r SIM-o
SMLLE(Sun et al., 2025) 5.14 6.37 0.516 0.489
IST-LM(Yang et al., 2024) - 4.53 - 0.653
BELLE-stream 3.54 4.44 0.524 0.491
We evaluate BELLE in a streaming setup to demonstrate its practical superiority in modern conversational pipelines. While we do not propose a novel streaming architecture, this evaluation serves a critical purpose: it validates that our Bayesian evidential framework is orthogonal to and fully compatible with standard autoregressive streaming. Unlike diffusion-based TTS models, which typically suffer from high algorithmic latency due to iterative denoising, BELLE retains the ability to synthesize “on-the-fly” without compromising the generation quality. This confirms that our approach effectively solves the trade-off between high-quality probabilistic modeling and the low Time-to-First-Audio (TTFA) required for duplex dialogue systems.
Table 5 presents the quantitative comparison. BELLE-stream achieves the lowest WER among the streaming systems (3.54%), significantly outperforming SMLLE and IST-LM. Crucially, there is no appreciable degradation in WER compared to the non-streaming BELLE, confirming that evidential modeling imposes no penalty on streaming performance.
In terms of efficiency, BELLE-stream achieves a real-time factor (RTF) of 0.55 for a 10-second utterance. With a chunk size of 0.8 seconds, this results in a first-packet latency (FPL) of approximately 440ms, demonstrating its suitability for real-time applications.
@@ 7 Conclusion
In this work, we introduced BELLE, a continuous autoregressive TTS framework that marks a fundamental shift from deterministic point estimation to principled Bayesian inference. By replacing fixed-variance priors with a dynamic Normal-Inverse-Gamma distribution, BELLE explicitly models the intrinsic “one-to-many” aleatoric uncertainty of human speech. To enable robust variance estimation on standard single-reference datasets, we proposed a novel training strategy that leverages synthetic samples not as imitation targets, but as a statistical support set to characterize the solution space.
Empirical results demonstrate that this evidential approach yields significant gains in robustness and naturalness: BELLE trained on $\sim$5k hours of data outperforms leading open-source models trained on 50k hours, achieving a 25.8% relative reduction in WER and superior speaker similarity. Crucially, these improvements are achieved without adding model parameters or inference latency, and the framework naturally extends to high-quality streaming scenarios.
Impact Statement
@@ Appendix A Limitations
Although BELLE demonstrates strong performance, we acknowledge several limitations in the current study.
First, the current work focuses solely on English speech data without validating multilingual generalization. Extending BELLE to diverse languages and verifying its cross-lingual capabilities remains an important direction for future work.
Second, regarding the experimental setup, we note that the official implementation of MELLE is not publicly available. Consequently, we rely on a faithful re-implementation as the baseline. While potential discrepancies in preprocessing pipelines (e.g., phoneme alignment) may cause the absolute performance of our MELLE baseline to differ slightly from the original paper, the comparison remains rigorously fair and valid, as both BELLE and the MELLE baseline utilize an identical training and evaluation pipeline in our experiments.
Third, due to computational resource constraints, the ablation studies (Table 4 and Table 8) were conducted on a reduced-scale dataset. While the absolute metric values in these studies are not directly comparable to the full-scale results (Table 2), this standard academic practice effectively isolates and verifies the relative contributions of individual components within the proposed framework.
Fourth, we acknowledge that our training data scale (under 5,000 hours) is considerably smaller than that of recent large-scale systems such as F5-TTS and MaskGCT (approx. 50,000 hours). However, this constraint effectively highlights a key strength of BELLE: it achieves comparable or even superior performance to these data-intensive baselines using an order of magnitude less data, demonstrating the exceptional data efficiency of our Bayesian evidential framework.
Finally, as with most autoregressive models, BELLE’s Real-Time Factor (RTF) is higher than that of recent non-autoregressive approaches. Further research on predicting multiple mel-spectrogram frames per step could substantially reduce RTF and enable even lower-latency streaming synthesis.
Appendix B Theoretical Clarification on the Bayesian Framework
In this section, we provide a deeper theoretical discussion regarding the classification of BELLE within the Bayesian Deep Learning landscape. While some strict definitions of Bayesian Neural Networks (BNNs) focus exclusively on placing priors over network weights (weight uncertainty), BELLE follows the Deep Evidential Learning (EDL) paradigm (Amini et al., 2020), which is mathematically grounded in Empirical Bayes (or Type-II Maximum Likelihood). We clarify the rigorous Bayesian nature of our approach and its utility in three aspects.
B.1 Empirical Bayes and Type-II Maximum Likelihood
BELLE does not treat the network weights as random variables, but instead treats the distributional parameters of the data likelihood as random variables. By placing a conjugate Normal-Inverse-Gamma (NIG) prior over the Gaussian likelihood parameters $(\mu,\sigma^{2})$, we formulate the training objective as maximizing the marginal likelihood (or model evidence):
$\log p(\bm{y}\mid\bm{x})=\log\int p(\bm{y}\mid\bm{\theta})p(\bm{\theta}\mid\bm{x})\,d\bm{\theta},$ (16)
where $\bm{\theta}=(\mu,\sigma^{2})$. This process analytically marginalizes over the aleatoric uncertainty parameters, effectively performing Type-II Maximum Likelihood Estimation. This distinguishes BELLE from standard regression (which optimizes point estimates) and firmly positions it within the Bayesian framework of evidence optimization.
B.2 Posterior Sampling in Generative Inference
Contrary to the view that the model “lacks posterior sampling,” BELLE explicitly performs sampling from the predicted posterior during the inference phase. Specifically, the network predicts the hyperparameters $\bm{m}=(\bm{\gamma},\bm{\nu},\bm{\alpha},\bm{\beta})$ of the NIG posterior distribution. The generative process then involves a hierarchical sampling procedure (as defined in Eq. 10):
  1. 1.
    Variance Sampling: We first sample the variance $\bm{\sigma}^{2}$ from the Inverse-Gamma posterior $\Gamma^{-1}(\bm{\alpha},\bm{\beta})$.
  2. 2.
    Mean Sampling: Conditioned on this variance, we sample the mean $\bm{\mu}$ from the Normal distribution $\mathcal{N}(\bm{\gamma},\sigma^{2}/\nu)$.
This stochastic sampling is critical: it breaks the deterministic “one-to-one” mapping typical of standard regression models, allowing BELLE to model the intrinsic “one-to-many” nature of human speech generation.
B.3 The Utility of Uncertainty: Addressing Oversmoothing
Finally, the learned uncertainty serves a crucial role in decision-making within the context of generative modeling. Standard regression approaches (based on MSE) tend to output the “average” of all possible prosodies to minimize error, leading to the well-known “oversmoothing” effect in TTS. In contrast, by explicitly modeling the higher-order uncertainty, BELLE identifies regions of high acoustic ambiguity (e.g., rapid transitions or expressive intonations) and dynamically allocates probability mass. This allows the model to sample valid, distinct acoustic realizations rather than collapsing to the mean, directly resulting in the enhanced prosodic diversity and naturalness observed in our experiments.
B.4 Theoretical Comparison with Standard Gaussian Sampling
Beyond the classification of the framework, we provide a rigorous theoretical comparison demonstrating why the proposed Evidential Deep Learning (EDL) approach is superior to the standard Gaussian sampling used in baseline models like MELLE.
Expressive Capacity (Student-$t$ as a Superset)
Theoretically, the EDL framework offers strictly greater expressive capacity. By marginalizing over the conjugate Normal-Inverse-Gamma (NIG) prior, the posterior predictive distribution follows a Student-$t$ distribution (Amini et al., 2020). Since the Gaussian distribution is mathematically defined as the limiting case of the Student-$t$ distribution as the degrees of freedom (represented by the evidence parameter $\nu$) approach infinity, our model’s capacity inherently subsumes that of the standard Gaussian approach. This ensures that BELLE retains the ability to model standard Gaussian distributions while providing the crucial additional flexibility to capture heavy-tailed distributions and complex outliers in the mel-spectrogram space.
Dynamic Variance vs. Fixed Regularization
A fundamental limitation of MELLE’s Gaussian sampling lies in its training objective. As shown in Eq. (4.2), MELLE minimizes the KL divergence against a prior distribution with a fixed unit variance ($\mathcal{N}(\bm{y}_{t},\bm{I})$). This term acts as a regularizer that forces the predicted variance towards $\bm{I}$, implicitly imposing a homoscedastic assumption (i.e., treating uncertainty as uniform across all speech frames). In contrast, our Bayesian approach enables dynamic variance estimation. By utilizing Type-II Maximum Likelihood, the model learns to estimate aleatoric uncertainty directly from the input evidence without the inductive bias of a fixed prior. This allows BELLE to effectively model the inherent heteroscedasticity of speech—dynamically assigning higher variance to ambiguous pronunciations and lower variance to deterministic segments—thereby improving robustness and generation quality.
Appendix C Modeling Temporal and Spectral Dependencies
A potential theoretical concern regarding the proposed framework is the assumption of a diagonal covariance matrix in the output distribution (i.e., independent Gaussian/NIG distributions per log-mel bin). Critics might argue that this factorization neglects the complex temporal and cross-band (spectral) correlations inherent in physical speech signals. In this section, we clarify how BELLE effectively captures these dependencies through its autoregressive backbone.
While the output distribution parameters at timestep $t$ are factorized dimension-wise, the hidden representations from which they are derived are deeply contextual and correlated. The model does not predict parameters based on an isolated frame; rather, the prediction is conditioned on a high-dimensional hidden state $\bm{h}_{t}$ generated by the Transformer decoder.
Mathematically, let $\bm{\Phi}_{t}=[\bm{\gamma}_{t},\bm{\nu}_{t},\bm{\alpha}_{t},\bm{\beta}_{t}]$ denote the predicted NIG distribution parameters at time $t$. The generation process is governed by:
$\bm{\Phi}_{t}=f_{\text{proj}}(\bm{h}_{t}),$ (17)
where $f_{\text{proj}}$ is the linear projection head. Crucially, the hidden state $\bm{h}_{t}$ is computed autoregressively:
$\bm{h}_{t}=\text{TransformerBlock}(\bm{e}_{<t},\bm{h}_{<t}),$ (18)
where $\bm{e}_{<t}$ represents the history of acoustic tokens.
Therefore, although the projection $f_{\text{proj}}$ maps $\bm{h}_{t}$ to factorized parameters, $\bm{h}_{t}$ itself serves as a compact summary of the entire temporal history and spectral context.
  • Temporal Dependency: Since $\bm{h}_{t}$ encapsulates information from $\bm{h}_{1:t-1}$, the predicted parameters $\bm{\Phi}_{t}$ are strictly conditioned on the temporal trajectory, ensuring physical continuity and coherence over time.
  • Spectral Dependency: The hidden state $\bm{h}_{t}$ is a dense vector where all frequency bands effectively interact within the self-attention mechanism layers before the final projection. Thus, cross-band correlations are implicitly modeled within the latent space before being disentangled at the final output layer.
Consequently, the factorization at the output layer is a design choice for computational tractability and closed-form sampling, which does not compromise the model’s ability to capture complex dependencies encoded in the autoregressive hidden states.
Appendix D Detailed Experimental Setup
D.1 Training Data